OpenStudy (anonymous):

A cylindrical trash can is 4 feet tall and holds 28.26 cubic feet of trash. What is the area of its base?

OpenStudy (anonymous):

it a cylindrical Volume of a cylinder = base * h

OpenStudy (anonymous):

28.26=4 * base base= 28.26/4=7.065